OpenStudy (anonymous):

Evaluate the expression when x = 4. 3x + 2 A. 36 B. 24 C. 18 D. 14

OpenStudy (anonymous):

D

OpenStudy (anonymous):

3(4)+2 3*4= 12 12+2=?

OpenStudy (uri):

@nikkittyuoy Put the value of x as 4 SO, 3(4)+2 12+2 =14 ans.

OpenStudy (anonymous):

ok